Rapid City traded in a used pickup for a new pickup with a sticker price of $44,000. The old truck had a historical cost of $40,000, accumulated depreciation of $16,000, and a fair value of $27,000. The dealer took the old truck and $13,000 cash for the new truck.

14. What is the amount of gain/loss that should be reported in the government-wide financial statements?
a) $0.
b) $3,000 gain.
c) $3,000 loss.
d) $7,000 gain

Need an explanation and formula
Answer #1

Total cost of truck should be reported = Fair value of old pickup + Cash for new truck

= 27,000+13,000

= 40,000

Calculation of book value of old truck:

Book value of old pickup = Historical cost of the old truck - Accumulated depreciation

= 40,000-16,000

24,000

Calculation of gain/loss

Loss on transaction = Book value of old pickup - Fair value on the date of transaction

= 24,000 - 27,000

= -3,000

So the amount of gain should be recognized on this transaction in the government-wide financial statement is 3,000.